OpenStudy (anonymous):

derive the equation of the parabola with a focus at (0.4) and a directix of y = 4

OpenStudy (anonymous):

@Hero Heyyy! . Are you familiar with this?

OpenStudy (anonymous):

@hartnn Hey are you familiar with this?

OpenStudy (anonymous):

the focus is at x=0, y=4 and the directrix is y=4?

OpenStudy (anonymous):

Yes correct (0,4)

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

this one is confusing for me, i'm not sure what to do when the directrix and focus have the same y-value. as far as i can figure, it's not a parabola, but it's a line...

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

Ugh ...math..

OpenStudy (anonymous):

Ive looked at help videos and i got so lost

OpenStudy (anonymous):

there are 3 equations that relate focus and diretrix to the equation of a parabola Ax^2+Bx+c focus x=-B/(2A) y=c-(b^2-1)/4a directrix y=c-(b^2+1)4a .. but the two last equations will both equal 4, which is what i'm stuck at becuase one equation has a +1 in it while the other has a -1, so i'm not sure...

OpenStudy (anonymous):

The one I learned was 1/4(p) = a

OpenStudy (anonymous):

and to find the focal length and vertex

OpenStudy (anonymous):

1/4(p) = a this is a good eq. but if we plug in 0 for 'p' we get a= undefined.... so thats why I don't think it's a parabola, but it is a line

OpenStudy (anonymous):

Im gonna call my teacher ...thanks :)

OpenStudy (agent0smith):

I'd guess it's a mistake.... you can't have the focus at the same spot as the directrix. Is it (0, 4) for the focus? and a directix of y = 4? no negatives?

OpenStudy (anonymous):

Yes @agent0smith